How many ounces of pure chocolate must be added to 140 oz of chocolate topping that is 50% chocolate to make a topping that is 75% chocolate?
? oz

Answer by stanbon(75887)

How many ounces of pure chocolate must be added to 140 oz of chocolate topping that is 50% chocolate to make a topping that is 75% chocolate?
----------------
Equation:
chocolate + chocolate = chocolate
0.50*140 + 1.00x = 0.75(140+x)
---
Multiply thru by 100 and solve for "x":
50*140 + 100x = 75*140 + 75x
25x = 25*140
x = 140 oz (amt. of pure chocolate that must be added to the mixture)
====================
Cheers,
Stan H.